OpenStudy (anonymous):

if six teachers can mark 36 papers in an haur and a half, how many papers can eight teachers mark in an hour? if six teachers can mark 36 papers in an haur and a half, how many papers can eight teachers mark in an hour? @Mathematics

OpenStudy (anonymous):

how do i set it up

OpenStudy (lilg132):

ok let me explain it properly

OpenStudy (anonymous):

ok

OpenStudy (lilg132):

u have 6 teachers that mark 36 papers in 90 minutes so first we work out how many papers one teacher can mark in one hour or 60 minutes. what we do is: 36 / 90 = 0.4 per minute now to get the answer for one hour we multiply by 60 0.4 * 60 = 24 but we need to work out how much per teacher so we divide that by 6 24 / 6 = 4 papers per teacher per hour now we need to work out how many papers would get marked with 8 teachers so you multiply that by 8 4 * 8 = 32 papers in one hour (oops my bad good u told me to show working out first answer was incorrect done in a hurry :p) so the answer is 32 papers with 8 teachers in one hour

OpenStudy (anonymous):

thank u very much!!!!!!!